Prepare 100mL of a 0.2M acetate buffer solution of pH =4.76 (using 17.6N acetic acid and sodium acetate).

a. Calculate the volume of acetic acid and weight of sodium acetate necessary to prepare the buffer solution. pKa of acetic acid is 4.76
Answer #1

Answer

Volume of 17.6N acetic acid required = 0.568ml

mass of sodium acetate required = 0.8204g

Explanation

mass of

Henderson-Hasselbalch equation is

pH = pKa + log([A-]/[HA])

4.76 = 4.76 + log( [CH3COO-] / [CH3COOH])

log([CH3COO-]/[CH3COOH]) = 0

[CH3COO-]/[CH3COOH] = 1

[CH3COO-]= [CH3COOH]

buffer concentration = 0.2M

[CH3COOH] = 0.1M

[CH3COO-] = 0.1M

Volume of CH3COOH required = 0.1M × 100ml/ 17.6M = 0.568ml

moles of CH3COO- required =( 0.1/1000ml)× 100ml = 0.01mol

moles of CH3COONa required = 0.01mol

mass of CH3COONa required = 0.01mol × 82.04g/mol

mass of CH3COONa required = 0.8204g